OpenStudy (anonymous):

please help with 2 questions! will give medal and become fan!

OpenStudy (anonymous):

1. Heather leaves on a long trip driving at a steady rate of 40 miles per hour. her sister Megan leaves from the same location traveling to the same destination 3 hours later. she drives at a steady rate of 50 miles per hour. how long after Megan leaves home will she catch up to heather? 2. you drop a rock off a bridge. the rocks height h (in feet above the water), after t seconds is modeled by h = -16t2 + 568. what is the height of the rock after 4 seconds

OpenStudy (neer2890):

for 2nd question put t=4 in your equation so h=-16*16+568=312 feet

OpenStudy (anonymous):

do you know 1?

OpenStudy (neer2890):

 let us suppose A to B is the distance travelled by heather in 3 hours. so Distance = speed*time AB=40*3=120 miles additionally, suppose BC is the distance travelled by both heather and megan and both meet together at point C.

OpenStudy (neer2890):

As heather and megan is going to travel the same distance upto point C. So,when distance is constant, we have \[\frac{ S1 }{ S2 }=\frac{ T1 }{ T2 }\] where S1,S2,T1,T2 are speeds and times taken by heather and megan. if we take left hand sideS1/S2 \[\frac{ 50 }{ 40 }=\frac{ 5 }{ 4 }\] which means \[\frac{ T1 }{ T2 }=\frac{ 5 }{ 4 }\]

OpenStudy (neer2890):

so it means when heather travels 4m by the same time megan travels 5m distance. so we can say that BC=4 and AC=5 which implies AB=1 (all these AB,AC and BC are in ratios,so if i'm saying AB=1 and BC=4 then this means AB:BC::1:4) we know, when AB=1 then BC=4 but AB=120 miles so when AB=120 then BC must be 120*4=480 miles then AC=480+120=600 miles. So, Megan will catch up to heather by 600 miles.